# Mix-of-Granularity: Optimize the Chunking Granularity for Retrieval-Augmented Generation

### v1

@@ Abstract
Integrating information from different reference data sources is a major challenge for Retrieval-Augmented Generation (RAG) systems because each knowledge source adopts a unique data structure and follows different conventions. Retrieving from multiple knowledge sources with one fixed strategy usually leads to under-exploitation of information. To mitigate this drawback, inspired by Mix-of-Expert, we introduce Mix-of-Granularity (MoG), a method that dynamically determines the optimal granularity of a knowledge database based on input queries using a router. The router is efficiently trained with a newly proposed loss function employing soft labels. We further extend MoG to Mix-of-Granularity-Graph (MoGG), where reference documents are pre-processed into graphs, enabling the retrieval of relevant information from distantly situated chunks. Extensive experiments demonstrate that both MoG and MoGG effectively predict optimal granularity levels, significantly enhancing the performance of the RAG system in downstream tasks. The code of both MoG and MoGG will be made public.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) [30] has become a popular method for enhancing Large Language Models (LLMs). The core concept of RAG involves retrieving relevant information from external knowledge bases to provide additional context to the LLM, enabling it to generate more precise and grounded responses. RAG offers a promising and practical solution to mitigate LLMs’ hallucinations because (1) it can be applied to any LLM, even those accessible only via APIs, and (2) the reference information is easy to modify or update. Many LLM-based products are supported by RAG systems, with examples spanning various industries such as customer service, advertising and marketing, education and e-learning, healthcare, and e-commerce and retailing [27, 51, 44, 26, 36].
@@
The quality and pertinence of the retrieved reference documents are crucial for the quality of the final generated response. Consequently, much research has focused on the retrieval phase. Currently, most RAG systems follow the Dual-Encoder Architecture [11] (DEA) paradigm, in which the reference documents are divided into small snippets, encoded by specific encoders, and then stored in the vector database (e.g. FAISS [24] or Neo4j [8]) as embeddings. The DEA paradigm shows great potential for empowering LLMs with any existing knowledge database (knowledge graphs, textbooks, Wikipedia, etc.), providing a practical solution to alleviate LLMs’ hallucinations.
However, a notable challenge in adopting the DEA paradigm is that the chunking size of reference documents is hard to optimize. For instance, effectively retrieving reference knowledge from medical textbooks or a Medical Knowledge Graph Database (e.g. Hetionet [16]) requires different chunking sizes because the data structure and information density in both sources are quite different: medical textbooks contain long text passages organized around similar topics, while a Medical Knowledge Graph Database is formed with entities of different terms (usually quite short) and the relations between them. In practice, different chunking sizes should be set for these two knowledge databases. However, searching for the optimal chunking size for each knowledge source is tedious. Another challenge of the DEA paradigm is that even when an optimal chunking size is determined, this fixed chunking size does not work well in all cases, because different types of users’ input correspond to different optimal chunking sizes. As illustrated in Figure 1, when the user asks for precise information about one disease or one medicine (fine-grained question), chunking the reference document in finer granularity is better; whereas, if the user asks for broad information, like the general information of a disease (coarse-grained question), a more coarse granularity of chunking is preferred. Thus, the community calls for a method to dynamically determine the chunking size, for which we propose the Mix-of-Granularity (MoG). We draw inspiration from Mix-of-Experts [6], which is a machine learning architecture that dynamically combines predictions from multiple specialized models, allowing for adaptive selection of the most relevant expertise for each input. Similarly, the fundamental concept of MoG involves training, in a supervised learning manner, a router to select and combine reference snippets of different granularity levels so that only the most pertinent snippets are passed to the backbone LLM. With the snippets of improved quality, the LLM can give better-grounded responses.
@@
Figure 1: The Mix-of-Granularity (MoG) mechanism can automatically select the optimal granularity when extracting information from the reference knowledge database (scenarios 1 and 2), achieving both high information pertinence and coverage. The selected text is passed to LLM for generation. When relevant information is dispersed across distant sections (scenario 3), MoG-Graph (MoGG) is applied to effectively retrieve these separate pieces of information. In MoGG, the reference documents are pre-processed in the form of a graph, and MoGG retrieves information with the best hopping range. The text corresponding to the selected subgraph is then fed to LLM for RAG.
@@
The MoG still has difficulty dealing with broader queries that require information from different documents of a single knowledge database or even information from different knowledge databases. In such cases, i.e., cross-document question in Figure 1, adjusting the granularity of the reference document is not helpful because the necessary information is so distantly located that it cannot be covered by a reasonably large chunking window. To better answer these broader queries, we extend MoG to Mix-of-Granularity-Graph (MoGG). In MoGG, the reference documents in the knowledge databases are pre-processed as a graph, allowing relevant snippets to be included as neighbors of each other, regardless of their distance in the original databases. This extension further improves the performance of retrieving information from multiple knowledge sources.
@@
In most existing RAG systems, when retrieving information from a database, a top-$k$ selection is implemented to control the number of candidate snippets. This common practice obstructs the backward propagation of the gradient, making the training infeasible. In this paper, we keep this top-$k$ selection for maximum compatibility with other works. To solve the problem of backward propagation, we introduce a loss function using soft labels. Soft labels are approximate training signals generated using offline algorithms or models like TF-IDF [38] or RoBERTa [32]. With the soft labels, the top-$k$ retrieval is excluded from the training process, thus the issue of backward propagation is circumvented and the training is accelerated.
In conclusion, the main contributions of this work are: (1) We propose MoG, which dynamically retrieves snippets from the optimal granularity level with the help of a router, achieving a balanced trade-off between coverage and pertinence during the retrieval phase of any RAG system. (2) We extend MoG to MoGG by reorganizing the reference document in the form of a graph, thereby further improving the quality of retrieval when relevant information is dispersed across distant sections in the knowledge database. (3) We introduce a loss function utilizing soft labels to overcome the challenges associated with training with top-$k$ selection.
@@ 2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) [30] has emerged as a standard practice to enhance the performance of LLMs, aiming to mitigate their problems of “hallucinations” and knowledge cut-off. An RAG system typically includes a Retriever that extracts relevant information from an external knowledge database, and a backbone LLM to generate grounded responses by considering the given relevant information (in-context learning [10]).
Technological advancements in RAG have addressed fundamental questions such as “what should be retrieved”, “when should retrieval occur”, and “how should the retrieved information be utilized”. In terms of “what to retrieve”, research has evolved from simple token retrieval [28] and entity retrieval [34] to more complex structures like chunks [37] and knowledge graphs [25]. Granularity matters a lot in retrieval, coarse-granularity-retrieval yields more information but with lower precision, while fine-granularity-retrieval offers comprehensive information at the cost of efficiency. The question of “when to retrieve” has led to strategies ranging from single [49, 42] to adaptive [20, 17] and multiple retrieval [19] methods. Regarding “how to use” the retrieved data, integration techniques have been developed for various levels of the model architecture, including the input [29], intermediate [3], and output layers [31]. Although integration at the intermediate and output layers is more effective, challenges remain concerning training requirements and efficiency limitations.
@@ 2.2 Chunking Optimization
Optimal chunk size is crucial for managing external documents. Breaking them down into small chunks is the first step to extracting and encoding the fine-grained features. Naive chunking strategies, such as “fixed-size” chunking and “recursive chunking,” attempt to create snippets of identical or similar size. However, fixed chunking size often yields suboptimal retrieval results in practice.
Current research in RAG explores chunking optimization techniques to improve retrieval efficiency and accuracy. One line of work focuses on increasing the recall of relevant snippets. For example, the “Sliding Window Chunking” [40] allows layered retrieval by merging globally related information across multiple processes. The “Parent Document Retrieval” [46] retrieves with small chunks and returns larger blocks of context for language model processing. Another line of work seeks to include more semantic information of the context to improve retrieval accuracy. “Metadata Filtering” [43] leverages document metadata to enhance the filtering process; “Context-Enriched Chunking” [45] breaks down information into meaningful segments and adds helpful summaries; while “Windowed Summarization Chunking” [45] enriches each text chunk with a windowed summary of the previous few chunks. Combining these methods has led to notable advancements, enhancing retrieval outcomes and improving performance in RAG.
@@ 2.3 Graph-Based Text Processing
Graph-based text processing techniques combine research in graphs and text retrieval. Previous works exploit semantic similarities between small snippets (a sentence or several words) and reorganize the text material into a graph using Entity Recognition and Relation Construction algorithms [33, 14, 52]. This approach breaks the constraint of the single dimension in the original text corpus, allowing chunks of the same topic to be grouped as neighbors in a graph. These graph-based text processing methods show great potential in tasks requiring reasoning over long texts or multi-hop reasoning.
Representative works in this field include “abstract embedding” [13], which prioritizes top-$k$ retrieval based on document abstracts (or summaries), providing a comprehensive understanding of the context. RAPTOR [41] organizes snippets as a tree (a special form of a graph) by recursively clustering them, where all non-leaf nodes correspond to summaries of their child nodes. This processing allows access to information at different granularity levels, resulting in significant improvements in summarization tasks. In GMoE [48], the authors use different expert networks to handle hop-$1$, hop-$2$, and mixed hop-$1$ hop-$2$ neighbors of a node in a graph, which inspires our extension from MoG to MoGG.
@@ 3.1 Preliminaries
Before describing the details of MoG and MoGG, let us begin with the basic annotations. MoG and MoGG are designed to enhance an existing RAG system, particularly its Retriever module. A typical RAG system comprises a Retriever $\mathcal{R}$, a Generator $\mathcal{G}$, and an external Knowledge database $\mathcal{K}$ (knowledge source). A RAG system takes a user query $q$ as input, retrieves reference document pieces (snippets or chunks) from $\mathcal{K}$, and uses these snippets to help the $\mathcal{G}$ produce the final responses.
In practice, the most popular architecture for the Retriever is the Dual-Encoders Architecture [11], where the query $q$ and all the snippets in $\mathcal{K}$ are encoded into embeddings ($e_{q}$ and $e_{s}$) using the same encoder $\mathcal{E}$. The extraction of relevant snippets is achieved by calculating the similarity between $e_{q}$ and $e_{s}$. The calculation of cosine similarities is made efficient thanks to modern vector databases like FAISS [24] or the Neo4j Database [8]. The Generator module is typically a backbone LLM, such as GPT [4], Llama [1], GLM [12], or InternLM [5]. Once the snippets are retrieved, they are usually injected into the backbone LLM via prompt injection. This method is favored because many powerful LLMs are accessible only via API calls. With the knowledge-integrated prompt, the LLM generates the final response. In some RAG systems, post-processing modules are added after generation, but this aspect is not covered in our work as it is not our focus.
@@ 3.2.1 Multi-granularity Router
We apply the idea of Mix-of-Expert [6] (MoE) to automatically determine the best granularity level in the retrieval phase of RAG. In a MoE system, different input tokens will be routed to the best expert network to be processed efficiently. Similarly, in MoG, a router is trained to predict the importance weight of different granularity levels based on the user’s input, so that the information retrieved from the optimal granularity level can be prioritized in the final top-$k$ selection and thus has a higher chance to be passed to the LLM.
Figure 2 illustrates the workflow of retrieving relevant snippets with a router. At the beginning, the user’s input query $q$ is encoded via RoBERTa before being mapped to a weight $w$ by the router. In parallel, the input query is used to retrieve relevant snippets from the knowledge database $\mathcal{K}$. In each $\mathcal{K}$, the documents within it are first chunked into $n_{gra}$ different granularity levels. Each snippet is then encoded into embeddings. The same retriever will retrieve the $k_{r}$ most relevant snippets (in Figure 2 $k_{r}$=3) from each granularity level based on their similarity with $q$. After the retrieval, a pool of candidate snippets of size $n_{gra}\times k_{r}$ is formed. During the retrieval, each candidate snippet is assigned a relevance score. Since all snippets are retrieved with the same retriever, their relevance scores are comparable. These scores are then weighted and summed using weight $w$ generated by the router. The $k$ snippets with the highest weighted relevance scores ($k$=1 in Figure 2) are selected and fed to the backbone LLM, helping LLM generate grounded responses.
Figure 2: Mix-of-Granularity (MoG) mechanism prioritizes the chunks retrieved from optimal granularity level, which is determined by the router based on the user input query.
@@
To trace the relevance scores of the same snippet across the different granularity levels, the documents are chunked in a non-overlapped way, and all chunks of coarser granularity are concatenations of several chunks of the finer granularity. In other words, each snippet of the finest granularity is included in one and only one snippet of coarser granularity. This approach allows us to trace the relevance scores of different granularity levels by the IDs of the finest chunks. For each chunk of the finest granularity level, if it is retrieved among the top-$k_{r}$ snippets from one granularity level, its relevance score for that level is recorded as the actual relevance score, just like the chunk marked in red in Figure 2; otherwise, it is filled with 0, like some zeroes padded for the chunk marked in blue. In this way, we get the relevance scores $t_{rs}$ of each finest snippet. Noting the set of all chunks of the finest granularity level as $\mathcal{C}$, the retrieved chunk ($chunk_{r}$) is selected by top-$k$ selection as follows:
$chunk_{r}=\underset{c\in\mathcal{C}}{\operatorname{top_{k}-argmax}}(t_{rs}(c)\cdot w).$ (1)
The snippet injected to the prompt of LLM is the snippet containing $chunk_{r}$ of the granularity level with the highest weight ($g_{r}$). In Figure 2, the snippet of the third granularity level is injected in the prompt because this level has the highest weight. Formally, $g_{r}$ is determined by:
$g_{r}=\underset{g\in[1,n_{gra}]|w_{g}\neq 0}{\operatorname{argmax}}w_{g},$ (2)
where $w_{g}$ represents the $g$-th element in $w$.

### v2

@@ Abstract
Integrating information from various reference databases is a major challenge for Retrieval-Augmented Generation (RAG) systems because each knowledge source adopts a unique data structure and follows different conventions. Retrieving from multiple knowledge sources with one fixed strategy usually leads to under-exploitation of information. To mitigate this drawback, inspired by Mix-of-Expert, we introduce Mix-of-Granularity (MoG), a method that dynamically determines the optimal granularity of a knowledge source based on input queries using a router. The router is efficiently trained with a newly proposed loss function employing soft labels. We further extend MoG to MoG-Graph (MoGG), where reference documents are pre-processed as graphs, enabling the retrieval of distantly situated snippets. Experiments demonstrate that MoG and MoGG effectively predict optimal granularity levels, significantly enhancing the performance of the RAG system in downstream tasks. The code of both MoG and MoGG are released in https://github.com/ZGChung/Mix-of-Granularity.
Mix-of-Granularity: Optimize the Chunking Granularity for Retrieval-Augmented Generation
Zijie Zhong Shanghai AI Laboratory zhongzijie@pjlab.org.cn                        Hanwen Liu Beihang University liuhwen@buaa.edu.cn                        Xiaoya Cui Beihang University xiaoya@buaa.edu.cn
Xiaofan Zhang [^1] Corresponding authors Shanghai AI Laboratory zhangxiaofan@pjlab.org.cn                        Zengchang Qin [^1] Corresponding authors Beihang University, theSight Technology zcqin@buaa.edu.cn
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) has become a popular method for enhancing Large Language Models (LLMs). The core concept of RAG involves retrieving relevant information from external knowledge bases to provide additional context to the LLM, enabling it to generate more precise and grounded responses. RAG offers a promising and practical solution to mitigate LLMs’ hallucinations because (1) it can be applied to any LLM, even those accessible only via APIs, and (2) the reference information is easy to modify or update. Many LLM-based products are supported by RAG systems, with examples spanning various industries such as customer service, advertising and marketing, education and e-learning, healthcare, and e-commerce and retailing Khan (2023); Xiong et al. (2024); Soman and Roychowdhury (2024); Ke et al. (2024); Radeva et al. (2024).
@@
The quality of the retrieved snippets is crucial for the final generation, consequently, much research has focused on the retrieval phase. Currently, most RAG systems follow the Dual-Encoder Architecture Dong et al. (2022) (DEA) paradigm, in which the reference documents are divided into small snippets (chunks), encoded by specific encoders, and then stored in the vector database (e.g. FAISS Johnson et al. (2017) or Neo4j Company (2012)) as embeddings. Thanks to its scalability, the DEA paradigm shows great potential for connecting LLMs with knowledge database of different formats, including knowledge graphs, textbooks, or Wikipedia articles.
@@
Figure 1: MoG automatically selects the optimal granularity when extracting information from the reference database (scenarios 1 and 2), achieving both high pertinence and coverage. When relevant information is dispersed across distant sections (scenario 3), the reference documents are pre-processed as graph, then MoGG is applied to retrieve these separate snippets from the best hopping range.
Optimizing the chunk size of the reference knowledge database is essential for enhancing the precision and recall during the retrieval phase. However, this optimization presents significant challenges in the implementation of RAG systems for two primary reasons: (1) An optimal chunking size needs to be determined for each knowledge database due to their dissimilar data structures and information densities. For example, the optimal chunking size for medical textbooks is longer than the one for Medical Knowledge Graph Database (MKGD like Hetionet Himmelstein et al. (2017)), as Medical textbooks typically contain lengthy passages, while an MKGD consists of entities represented by shorter terms. (2) Even within a single knowledge database, using a uniform chunking size for all input queries can yield suboptimal retrieval results, as the queries themselves exhibit varying levels of granularity. As shown in Figure 1), when the user asks about one disease (fine-grained question), chunking the reference document in finer granularity is better; whereas, when the user asks for broader information (coarse-grained question), a more coarse granularity is preferred. In practice, the optimal "uniform chunking size" is determined through parameter tuning, which is not only tedious but also does not guarantee high precision and recall.
Therefore, the community calls for a method to dynamically determine the optimal chunking size, for which we propose the Mix-of-Granularity (MoG). We draw inspiration from Mix-of-Experts Chen et al. (2022), which is a machine learning architecture that dynamically selects the most pertinent “expert sub-network” for each input token using a router module. Similarly, MoG involves leveraging a router to select the best granularity levels from which the reference snippets are extracted.
@@
Even with this flexibility newly introduced, MoG still has difficulty dealing with broader queries that require distantly situated snippets, e.g. snippets stored in different knowledge databases. In such cases (cross-document question in Figure 1), adjusting only the granularity is not helpful because the necessary information is so distantly located that it can never be covered by a reasonably large chunking window. To better answer these broader queries, we extend MoG to MoG-Graph (MoGG). In MoGG, the reference documents are pre-processed as a graph, allowing relevant snippets to be included as neighbors of each other, regardless of their distance in the original databases. This extension further improves the performance of cross-source retrieval.
@@
When training MoG(G) under the supervised learning setting, the backward propagation is blocked by the top-$k$ selection at the end of the retrieval phase, which is a common practice in most RAG systems. To solve it, we introduce a loss function using soft labels. Soft labels are approximate training signals generated using offline algorithms or models like TF-IDF Ramos (2003) or RoBERTa Liu et al. (2019). With the soft labels, the top-$k$ selection is excluded from the training process, thus the issue of backward propagation is circumvented and the training is accelerated.
In conclusion, the main contributions of this work are:
(1) We propose MoG, which dynamically determines the optimal granularity level for retrieval with the help of a router, achieving a balanced trade-off between precision and recall.
(2) We extend MoG to MoGG by reorganizing the reference document in the form of a graph, thereby further improving the quality of cross-source retrieval involving multiple knowledge databases.
(3) We introduce a loss function utilizing soft labels to overcome the challenges associated with training with top-$k$ selection.
@@ 2.1 Retrieval-Augmented Generation
RAG Lewis et al. (2020) has emerged as a standard practice to enhance the LLMs by mitigating their problems of “hallucinations” and knowledge cut-off. A RAG system typically includes a Retriever that extracts relevant information from an external knowledge database, and a backbone LLM to generate grounded responses by in-context learning Dong et al. (2023). Previous retrieval-focused methods have evolved from retrieving tokens Khandelwal et al. (2020) or entities Nishikawa et al. (2022) to more complex structures like chunks Ram et al. (2023) or graphs Kang et al. (2023). Granularity matters a lot in retrieval, as coarse-granularity-retrieval yields more information with lower precision, while fine-granularity-retrieval offers comprehensive information at the cost of efficiency. More strategies like single Wang et al. (2023b); Shi et al. (2023), adaptive Jiang et al. (2023); Huang et al. (2024), or multiple retrieval Izacard et al. (2022b) are introduced to improve the retrieval phase’s performance. Regarding the generation phase, various information fusion techniques are developed to integrate retrieval results to LLM in its input Khattab et al. (2023), intermediate Borgeaud et al. (2022), or output layers Liang et al. (2024).
@@ 2.2 Chunking Optimization
Optimal chunk size is crucial for RAG system as breaking down documents into small chunks is the first step to encode them. Naive chunking strategies, such as “fixed-size” chunking and “recursive chunking,” attempt to create snippets of identical size. Later works explored more chunking optimization techniques. One line of work focuses on increasing the recall of retrieval. For example, the “Sliding Window Chunking” Safjan (2023) allows layered retrieval by merging globally related information across multiple processes. The “Parent Document Retrieval” team (2023) retrieves using small chunks and returns larger blocks of context for later generation. Another line of work seeks to include more semantic information of the context to improve retrieval accuracy. “Metadata Filtering” Siegler (2024) leverages document metadata to filter snippets; “Context-Enriched Chunking” team (2024) breaks down information into segments and adds semantic summaries before retrieval; while “Windowed Summarization Chunking” team (2024) enriches each chunk with a windowed summary of the previous chunks. Many of these techniques, including MoG(G), are compatible with each other and can be combined to achieve better performance.
@@ 2.3 Graph-Based Text Processing
Graph-based text processing techniques combine research in graphs and text retrieval. Previous works exploit semantic similarities between small snippets (a sentence or several words) and reorganize the text material into a graph using Entity Recognition and Relation Construction algorithms Melnyk et al. (2022); Guo et al. (2020); You et al. (2018). Breaking the constraint of the single dimension of text corpus, these methods allow chunks of the same topic to be grouped as neighbors in a graph, thus show great potential in tasks requiring long context reasoning or multi-hop reasoning. For example, “graph indexing” Gao et al. (2024) transforms entities and relationships into nodes and connections, improving the relevance of retrieved snippets significantly. RAPTOR Sarthi et al. (2024) organizes snippets as a tree (a special form of a graph) by recursively clustering them, where all non-leaf nodes correspond to summaries of their child nodes. This processing allows access to information at different granularity levels, which is beneficial to summarization tasks. In GMoE Wang et al. (2023a), authors use different expert networks to handle hop-$1$, hop-$2$, and mixed hop-$1$ hop-$2$ neighbors of a node in a graph, which inspired our design of MoGG.
@@ 3.1 Preliminaries
MoG(G) is designed to enhance the Retriever of a RAG system. A typical RAG system comprises a Retriever $\mathcal{R}$, a Generator $\mathcal{G}$, and a series of external reference corpus $\mathcal{K}=\{\mathcal{K}_{1},\mathcal{K}_{2},...,\mathcal{K}_{k}\}$. A RAG system takes a user query $q$ as input, retrieves reference document pieces (snippets or chunks) from $\mathcal{K}$, and uses these snippets to help the $\mathcal{G}$ produce the final responses. A popular architecture for the Retriever is DEA Dong et al. (2022), where the query $q$ and all the snippets in $\mathcal{K}$ are encoded into embeddings ($e_{q}$ and $e_{s}$) using the same encoder $\mathcal{E}$. The extraction of relevant snippets is achieved by calculating the similarity between $e_{q}$ and $e_{s}$. The snippets with highest similarity scores are extracted and injected into the backbone LLM via prompt.
@@ 3.2.1 Multi-granularity Router
We apply the idea of Mix-of-Expert Chen et al. (2022) (MoE) to automatically determine the best granularity level in the retrieval phase. In a MoE system, different input tokens are routed to the best expert network based on the weights output by the router. Similarly, in MoG, a router is trained to predict the importance weight of different granularity levels based on the user’s input, so that the snippets from the best granularity level are prioritized. By employing such a routing optimization method, we can effectively adjust the chunk size according to different scenarios. Take one corpus $\mathcal{K}$ as example, its documents are chunked in $n_{gra}$ candidate granularities before the retrieval (in Figure 2, $n_{gra}$ = 5). Although some methods can dynamically adjust chunks using well-trained models, we prioritize the most commonly used and simplest chunking method from the perspective of practical efficiency. The snippets are chunked without overlap, and each chunk in $j$ ($j\in[2,n_{gra}]$) granularity level is formed with 2 adjacent chunks in $j-1$ granularity level (granularity level 1 is the finest). Each chunk is assigned a similarity score using BM25 Robertson and Zaragoza (2009) with respect to user’s input query $q$. In each granularity, $k_{r}$ most relevant snippets are extracted ($k_{r}$ = 3 in Figure 2), forming a pool of candidate snippets of size $n_{gra}\times k_{r}$. In parallel, $q$ is encoded via RoBERTa before being mapped to a weight $w$ (a vector of the same length as $n_{gra}$) by the router. The chunks’ similarity scores are then weighted and summed with $w$.
Empirically, directly retrieving the top snippets across different granularities based on their weighted similarity scores will give biased results, because the scores of more coarsed granularity levels are systematically higher (more discussions in Appendix A). Therefore, we adopt another selection strategy involving two steps: (1) select the top-$k$ ($k$ = 1 in Figure 2) finest-grained chunk $chunk_{r}$; (2) retrieve the chunk containing $chunk_{r}$ from the optimal granularity level $g_{r}$. The selection process can be formalized as follows:
$chunk_{r}$ $=\underset{c\in\mathcal{C}}{\operatorname{top_{k}-argmax}}(t_{rs}(c)\cdot w);$ (1)
$g_{r}$ $=\underset{g\in[1,n_{gra}]|w_{g}\neq 0}{\operatorname{argmax}}w_{g}\;$ (2)
where $\mathcal{C}$ represents the set of all chunks of the finest granularity level, $t_{rs}(.)$ represents the relevance score of a chunk, and $w_{g}$ represents the $g$-th element in $w$.
@@
Two snippets are highlighted in Figure 2 as examples to help understanding. For each chunk of the finest granularity level, if it is retrieved among the top-$k_{r}$ snippets from one granularity level, its relevance score for that level is recorded as the actual relevance score (like the red chunk); otherwise, it is filled with 0 (like some zeroes padded for the blue chunk). In this way, we get the relevance scores $t_{rs}$ of each finest snippet. The blue chunk has highest weighted similarity score, it is therefore selected as $chunk_{r}$. In this example, the final snippet selected is the one extracted from 3rd granularity level containing $chunk_{r}$.
Figure 2: MoG mechanism prioritizes the chunks retrieved from optimal granularity level, which is determined by the router based on the user input query.